# On the finite basis of two-loop ‘t Hooft-Veltman Feynman integrals

### v1

@@ 3 Evanescent sectors
In what follows, we will apply a similar reasoning at two loops. We start by reiterating that any higher-point two-loop integrals can be decomposed into subsectors of the 12 two-loop finite basis topologies Bargiela:2024rul , as depicted in the last 12 rows of Tab. 1. We will analyze all their subsectors with more than 8 denominators, i.e. 9, 10, and 11. We expect all of them to be evanescent since there are only $4L$ linearly independent propagators at $L$ loops in exactly $d=4$ dimensions. As such, this effect starts appearing for integrals with the number of denominators $n+3(L-1)$ exceeding $4L$, i.e. with at least $n=L+4$ legs. We start by running, the IBP reduction using Kira Maierhofer:2017gsa ; Klappert:2020nbg ; Klappert:2019emp for each of the 12 finite basis topologies while dropping subsectors with 8 or less denominators. To make the reduction feasible while preserving the representative properties of the IBP system, we select a generic numerical probe for the external kinematics and focus on all massless propagators. Specifically, we place the four basis momenta $p_{1},p_{2},p_{3},p_{4}$ on-shell and the other legs $p_{i>4}$ off-shell, and retain the full parametric dependence on $d$. Note that we analyze here also these subsectors which may have the same graph but differ in kinematic parametrization. We obtained $\{{167,167,167,295,285,287,332,327,332,445,429,445}\}$ MIs without any dots for the last 12 topologies in Tab. 1, respectively, while still dropping subsectors with 8 or less denominators.
@@
This concludes the construction of an evanescent MI basis for sectors with more than 8 denominators, on a numerical probe. Interestingly, we found that, when considering the IBP reduction in each sector separately, there are 7 sectors, as depicted in Fig. 1, each with 10 denominators, which are doubly evanescent, i.e. $M_{i}^{(\text{4D})}=\epsilon^{2}\,A^{\text{(fin)}}_{ij}M^{(\text{6D})}_{j}+\text{subsectors}$. We point out that, since our analysis was performed on a numerical kinematic sample, it is not a formal proof. Indeed, it would require solving the IBP system analytically for all the kinematic parameters in every possible degenerate limit separately, which is prohibitive at the moment. Recently, for specific topologies and kinematics, such an analytic proof has been provided using other methods in Ref. Abreu:2024fei for fully massless on-shell six-point double-pentagon and hexa-box CDR sectors. It would be interesting to confirm our numerical observation fully analytically, including the existence of the doubly-evanescent sectors.
@@
where $a_{i}$ and $b_{j}$ are rational functions of kinematic invariants, and they arise from the IBP reduction. Thus, we have to argue that $b_{j}$ are finite in $\epsilon$. This can be shown by assuming the usual property of all known $L$-loop amplitudes that their finite part has at most transcendental weight $2L$. Indeed, then $\epsilon$-divergent $b_{j}$ is excluded because, after multiplying with $M^{(2)}_{>8\text{den},\text{evan},j}\sim\mathcal{O}(\epsilon)$, there could be contributions to the finite part of weight higher than four. Such higher-weight terms would have to cancel for any generic process, i.e. arbitrary $b_{j}$ coefficients. This concludes the argument that all the two-loop Master Integrals, expressed in an appropriate basis, with more than 8 denominators, i.e. 9, 10, or 11, would not contribute to the finite part of any two-loop amplitude.
@@ 4.2 Partial fraction for DCT
Recently, a method called the dimension-changing transformation (DCT) Huang:2024qan has emerged as an effective approach for the efficient numerical evaluation of one-loop multi-point integrals to higher orders in $\epsilon$. In this section, we show that combining the closed partial fraction form proposed by us in Ref. Bargiela:2024rul with the DCT method — which we refer to as ”PF + DCT” — can further enhance the efficiency of numerical evaluation compared to evaluating CDR integrals using the DCT method, referred to as ”CDR + DCT.”
@@ 6 Conclusions
Our analysis induces natural steps to pursue in the future. Firstly, a lot of the finite basis MIs are still awaiting to be calculated explicitly. For the phenomenologically motivated kinematics, there has been a major effort in this direction, for two-loop four-point Chaubey:2024ckh , five-point Chicherin:2020oor ; Chicherin:2021dyp ; Badger:2024fgb , and six-point Abreu:2024fei ; Henn:2025xrc scattering. Secondly, it would be important to verify analytically that all of the two-loop MIs with more than 8 denominators can be chosen to be evanescent. For the massless on-shell double-pentagon and hexa-box, it has been already shown in Ref. Abreu:2024fei . Thirdly, the tHV decomposition will improve the performance of the IBP reduction for both integrals and amplitudes, and the associated numerical evaluation, in new high-multiplicity processes. Indeed, it has been already used for reducing the complexity of the numerical evaluation of the massless on-shell double-pentagon MIs in Ref. Henn:2025xrc . Finally, the classification of the special functions and the corresponding geometries would help in constructing a more efficient functional ansatzes for Feynman integrals and scattering amplitudes in the bootstrap approach Dixon:2011pw ; Chicherin:2017dob ; Hannesdottir:2024hke . In addition, it would be beneficial to further constrain the spectrum, e.g. by simplifying some of the hyper-surfaces stemming from the 30 corresponding graphs in Figs 5 and 7.
@@ Acknowledgements
We thank T. Gehrmann and R. Marzucca for valuable comments on this paper draft. We are grateful to G. Fontana, H. Frellesvig, F. Lange, R. Morales, S. Pögel, V. Sotnikov, M. Wilhelm, and S. Zoia for interesting discussions. This research was supported by the Swiss National Science Foundation (SNF) under contract 200020-204200 and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme grant agreement 101019620 (ERC Advanced Grant TOPUP).

### v2

@@ 3 Evanescent sectors
In what follows, we will apply a similar reasoning at two loops. We start by reiterating that any higher-point two-loop integrals can be decomposed into subsectors of the 12 two-loop finite basis topologies Bargiela:2024rul , as depicted in the last 12 rows of Tab. 1. We will analyze all their subsectors with more than 8 denominators, i.e. 9, 10, and 11. We expect all of them to be evanescent since there are only $4L$ linearly independent propagators at $L$ loops in exactly $d=4$ dimensions. As such, this effect starts appearing for integrals with the number of denominators $n+3(L-1)$ exceeding $4L$, i.e. with at least $n=L+4$ legs. We start by running the IBP reduction using Kira Maierhofer:2017gsa ; Klappert:2020nbg ; Klappert:2019emp for each of the 12 finite basis topologies while dropping subsectors with 8 or less denominators. To make the reduction feasible while preserving the representative properties of the IBP system, we select a generic numerical probe for the external kinematics and focus on all massless propagators. Specifically, we consider the four basis momenta $p_{1},p_{2},p_{3},p_{4}$ on-shell and the other legs $p_{i>4}$ off-shell, and retain the full parametric dependence on $d$. Note that the integrals in this kinematics are in general singular in $\epsilon$. In addition, we point out that we analyze here also these subsectors which may have the same graph but differ in kinematic parametrization. We obtained $\{{167,167,167,295,285,287,332,327,332,445,429,445}\}$ MIs without any dots for the last 12 topologies in Tab. 1, respectively, while still dropping subsectors with 8 or less denominators.
@@
This concludes the construction of an evanescent MI basis for sectors with more than 8 denominators, on a numerical probe. Interestingly, we found that, when considering the IBP reduction in each sector separately, there are 7 sectors, as depicted in Fig. 1, each with 10 denominators, which are doubly evanescent, i.e. $M_{i}^{(\text{4D})}=\epsilon^{2}\,A^{\text{(fin)}}_{ij}M^{(\text{6D})}_{j}+\text{subsectors}$. We point out that, since our analysis was performed on a numerical kinematic sample, it is not a formal proof. Indeed, it would require solving the IBP system analytically for all the kinematic parameters in every possible degenerate limit separately, which is prohibitive at the moment. Recently, for specific topologies and kinematics, the evanescence has been shown using other methods in Ref. Abreu:2024fei for fully massless on-shell six-point double-pentagon and hexa-box CDR sectors. It would be interesting to confirm our numerical observation fully analytically, including the existence of the doubly-evanescent sectors.
@@
where $a_{i}$ and $b_{j}$ are rational functions of kinematic invariants. They arise from the amplitude integrand in a finite manner, and from the IBP reduction which may induce poles in $\epsilon$. Thus, we have to argue that $b_{j}$ are finite in $\epsilon$. Similarly as in Ref. Abreu:2024fei , this can be shown by considering all the possible numerators allowed in renormalizable theories. The relevant maximal total numerator degree is bounded by the occurrence of momenta in numerators of Feynman rules applied to the particle dressings of the two-loop finite basis topologies. By IBP reducing all of the allowed integrals using Kira onto the evanescent MIs basis constructed above, we observed that the IBP coefficients are finite in $\epsilon$. Beyond 11 denominators, for each additional denominator, one more numerator power is allowed. Importantly, due to the linear dependence of the additional denominator on the basis 11 denominators, the one additional numerator can be always canceled, e.g.
$\begin{split}\frac{1}{\mathcal{D}_{1}\cdots\mathcal{D}_{11}}\frac{\mathcal{N}}{\mathcal{D}_{12}}=\frac{1}{\mathcal{D}_{12}}\frac{a_{0}+\sum_{i=1}^{11}a_{i}\mathcal{D}_{i}}{\mathcal{D}_{1}\cdots\mathcal{D}_{11}}=\frac{1}{\mathcal{D}_{12}}\bigg{(}\frac{a_{0}}{\mathcal{D}_{1}\cdots\mathcal{D}_{11}}+\sum_{i=1}^{11}\frac{a_{i}}{\mathcal{D}_{1}\cdots\hat{\mathcal{D}}_{i}\cdots\mathcal{D}_{11}}\bigg{)}\,,\end{split}$ (7)
where $\hat{\mathcal{D}}_{i}$ denotes an absence of the denominator $\mathcal{D}_{i}$. We point out that this numerator reduction is also important for decreasing the associated computational complexity. Finally, the remaining linearly dependent denominators can be partial fractioned into the subsectors of the finite basis topologies, according to Ref. Bargiela:2024rul . This concludes the argument that all the two-loop Master Integrals, expressed in an appropriate basis, with more than 8 denominators, i.e. 9, 10, or 11, would not contribute to the finite part of any two-loop amplitude.
@@ 4.2 Partial fraction for DCT
Recently, a method called the dimension-changing transformation (DCT) Huang:2024qan has emerged as an effective approach for the efficient numerical evaluation of one-loop multi-point integrals to higher orders in $\epsilon$. In this section, we show that combining the closed partial fraction form proposed by us in Ref. Bargiela:2024rul with the DCT method — which we refer to as ”PF + DCT” — can further enhance the efficiency of numerical evaluation compared to evaluating CDR integrals using the DCT method, referred to as ”CDR + DCT”.
@@ 6 Conclusions
Our analysis induces natural steps to pursue in the future. Firstly, a lot of the finite basis MIs are still awaiting to be calculated explicitly. For the phenomenologically motivated kinematics, there has been a major effort in this direction, for two-loop four-point Chaubey:2024ckh , five-point Chicherin:2020oor ; Chicherin:2021dyp ; Badger:2024fgb ; FebresCordero:2023pww , and six-point Abreu:2024fei ; Henn:2025xrc scattering. Secondly, it would be important to verify analytically that all of the two-loop MIs with more than 8 denominators can be chosen to be evanescent. For the massless on-shell double-pentagon and hexa-box, it has been already shown in Ref. Abreu:2024fei . Thirdly, the tHV decomposition will improve the performance of the IBP reduction for both integrals and amplitudes, and the associated numerical evaluation, in new high-multiplicity processes. Indeed, it has been already used for reducing the complexity of the numerical evaluation of the massless on-shell double-pentagon MIs in Ref. Henn:2025xrc . Finally, the classification of the special functions and the corresponding geometries would help in constructing a more efficient functional ansatzes for Feynman integrals and scattering amplitudes in the bootstrap approach Dixon:2011pw ; Chicherin:2017dob ; Hannesdottir:2024hke . In addition, it would be beneficial to further constrain the spectrum, e.g. by simplifying some of the hyper-surfaces stemming from the 30 corresponding graphs in Figs 5 and 7.
@@ Acknowledgements
We thank R. Marzucca for valuable comments on this paper draft. We are grateful to K. Schönwald, V. Sotnikov, and S. Zoia for important comments on the evanescent integrals. We also thank G. Fontana, H. Frellesvig, F. Lange, R. Morales, S. Pögel, and M. Wilhelm for interesting discussions. This research was supported by the Swiss National Science Foundation (SNF) under contract 200020-204200 and by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme grant agreement 101019620 (ERC Advanced Grant TOPUP).